# Is A Picture Worth A Thousand Words? Delving Into Spatial Reasoning for Vision Language Models

### v1

@@ 6 Discussion and Conclusions
We explored the spatial understanding capabilities of VLMs and LLMs. Our experiments resulted in several surprising conclusions across SoTA open-source and proprietary models. (1) VLMs struggle with spatial reasoning tasks (2) multimodal models rarely surpass LLMs when relying on visual inputs, (3) when both textual and visual inputs are provided, multimodal language models rely less on the visual inputs, and (4) VLMs often outperform their LLM counterparts with text-only inputs. This challenges the belief that current VLMs are highly performant at a wide range of vision-text tasks. However, with further thought, perhaps this is to be expected after all. The currently known architectures for VLMs attempt to “translate" the vision input into the language space and all reasoning is then performed in the language domain. It is logical that this automatic translation path is worse than a human provided translation to text, as in our text-only scenario. Thus our work shows the limits of the translation approach. Instead, to bridge the gap to human performance, future models require new architectures that treat vision input as a first-class source of information and reason in a joint vision-language space. It is our hope that our work informs development on this path.
@@
Appendix A Limitations and Societal Impact
@@
In this work, we introduce three novel benchmarks and conduct a comprehensive evaluation of diverse large language models (LLMs) and vision-language models (VLMs), presenting a controlled and thorough evaluation of spatial reasoning capabilities. While our analysis is detailed and extensive, it remains primarily empirical. We believe that embarking on a formal theoretical study, despite its challenges, would significantly enrich our understanding of pre-trained multimodal language models. Moreover, our focus has been on in-depth analysis rather than on developing new training strategies or adaptation algorithms to enhance spatial reasoning. Recognizing these points, we identify them as valuable directions for future research.
@@
Appendix B Experimental Details
B.1 Software and Hardware
Our experiments are conducted on 4 NVIDIA RTX A6000 GPUs (for small-scale models with fewer than 2B parameters) and 4 NVIDIA A100 GPUs (for 7B, 13B, and 34B models). Our implementation is based on Python 3.10 and PyTorch 2.1.2.
B.2 Hyperparameters and Error Bars
@@ B.3 Model Checkpoints
B.4 Detailed Illustration of Datasets
In the main text, we abbreviated the textual descriptions in Figure 1, Figure 2 due to space constraints. In this section, we provide the full descriptions for each task to facilitate a better understanding of spatial reasoning tasks. The complete illustrations are displayed in Figure 13 for Spatial-Map and Figure 14 for Maze-Nav, respectively.
@@ B.4 Detailed Illustration of Datasets
Appendix C Further Ablation Studies and Discussions
@@
Figure 15: TQA vs. VTQA on spatial reasoning tasks. Each vertex on the spider plot represents the Avg Acc of a (LLM, VLM) pair with the same language model backbone. For LLM, we use the Text-only input (i.e., TQA); for VLM, we use Vision-text input (i.e., VTQA).
Appendix D Detailed Experimental Results
In Section 4 and Section 5, to clearly illustrate the impact of modality, we present the results averaged over all questions in Figure 5 for VQA vs. VQA and Figure 10 for VQA vs. VTQA. This section provides a comprehensive breakdown of results for each individual question within the tasks. Detailed comparative results for Spatial-Map, Maze-Nav, and Spatial-Grid are shown in Table 3, Table 4, and Table 5, respectively. We compare language models (LM) and vision-language models (VLM) with the same language model backbone. For the VLM assessments, we consider inputs in both vision-only (VQA) and vision-text (VTQA) formats.
@@ Appendix D Detailed Experimental Results
Table 2: Detailed results for proprietary models
@@
Table 3: Detailed results for the Spatial-Map task. We compare language models (LM) and vision-language models (VLM) with the same language model backbone. For VLMs, we consider both vision-only and vision-text input format. The averaged results are summarized in Figure 5 and Figure 10.
@@
Table 4: Detailed results for the Maze-Nav task. We compare language models (LM) and vision-language models (VLM) with the same language model backbone. For VLMs, we consider both vision-only and vision-text input format. The averaged results are summarized in Figure 5 and Figure 10.
@@
Table 5: Detailed results for the Spatial-Grid task. We compare language models (LM) and vision-language models (VLM) with the same language model backbone. The averaged results are summarized in Figure 5 and Figure 10.

### v2

@@ 6 Discussion and Conclusions
We explored the spatial understanding capabilities of VLMs and LLMs. Our experiments resulted in several surprising conclusions across SoTA open-source and proprietary models. (1) VLMs struggle with spatial reasoning tasks, (2) multimodal models rarely surpass LLMs when relying on visual inputs, (3) when both textual and visual inputs are provided, multimodal language models rely less on the visual inputs, and (4) VLMs often outperform their LLM counterparts with text-only inputs. This challenges the belief that current VLMs are highly performant at a wide range of vision-text tasks. However, with further thought, perhaps this is to be expected after all. The currently known architectures for VLMs attempt to “translate" the vision input into the language space and all reasoning is then performed in the language domain. It is logical that this automatic translation path is worse than a human provided translation to text, as in our text-only scenario. Thus our work shows the limits of the translation approach. Instead, to bridge the gap to human performance, future models require new architectures that treat vision input as a first-class source of information and reason in a joint vision-language space. It is our hope that our work informs development on this path.
Acknowledgement
The authors would like to thank NeurIPS anonymous reviewers for their insightful feedback and helpful discussions. Yifei Ming and Yixuan Li are funded in part by the AFOSR Young Investigator Program under award number FA9550-23-1-0184, National Science Foundation (NSF) Award No. IIS-2237037 & IIS-2331669, and Office of Naval Research under grant number N00014-23-1-2643.
@@ Acknowledgement
Appendix A How Does SpatialEval Expand Beyond Traditional VQA?
We highlight several key rationales for SpatialEval in comparison to existing VQA benchmarks:
  • Task scope and focus: Existing benchmarks such as Visual Genome [31], GQA [25], and BLINK [18] focus only on VQA, where the image is required but the text description is often omitted or optional. In contrast, SpatialEval further explores spatial reasoning across different settings: TQA (LLM), TQA (VLM), and VTQA, where images or texts can be optional, thereby broadening the scope of tasks.
  • Evaluation scheme: we primarily focus on generative LLMs and VLMs which can “elaborate on the reasoning behind your answer in a detailed, step-by-step explanation” (Section 2.2). The task in prior VQA benchmarks is often treated as discriminative with no explicit reasoning. Therefore, it remains unknown if previous observations can be naturally transferred to foundation models pre-trained on web-scale data.
  • The Textual Representation of Images: The textual descriptions in prior VQA benchmarks are often brief or directly imply the answers. While these datasets are valuable, they do not consistently offer the level of complexity we require, such as numerous objects containing dense visual information along with detailed natural language captions that fully convey the image content. In SpatialEval, we provide long and dense captions for each image. As a result, answers cannot be easily inferred. We also aim to isolate object detection capability from spatial reasoning ability by simplifying objects to symbols (e.g., Spatial-Map).
  • IQ test for VLMs and LLMs: Tasks in SpatialEval are designed to serve as cognitive tests that evaluate basic capabilities of multi-modal foundation models. While three tasks feature synthetic visual content, humans can solve them with near-perfect accuracy. This indicates that the tasks are within the realm of human cognitive capabilities.
Appendix B Limitations and Societal Impact
@@
In this work, we introduce four novel tasks and conduct a comprehensive evaluation of diverse large language models (LLMs) and vision-language models (VLMs), presenting a controlled and thorough evaluation of spatial reasoning capabilities. While our analysis is detailed and extensive, it remains primarily empirical. We believe that embarking on a formal theoretical study, despite its challenges, would significantly enrich our understanding of pre-trained multimodal language models. Moreover, our focus has been on in-depth analysis rather than on developing new training strategies or adaptation algorithms to enhance spatial reasoning. Recognizing these points, we identify them as valuable directions for future research.
@@
Appendix C Experimental Details
C.1 Software and Hardware
Our experiments are conducted on NVIDIA A100 GPUs. Our implementation is based on Python 3.10 and PyTorch 2.1.2.
C.2 Hyperparameters and Error Bars
@@ C.3 Model Checkpoints
C.4 Detailed Illustration of Datasets
In the main text, we abbreviated the textual descriptions in Figure 1, Figure 2 due to space constraints. In this section, we provide the full descriptions for each task to facilitate a better understanding of spatial reasoning tasks. The complete illustrations are displayed in Figure 14 for Spatial-Map and Figure 15 for Maze-Nav, respectively. Three questions (termed Q1 to Q3) are associated with each sample in tasks Spaitial-Map, Maze-Nav, and Spatial-Grid.
@@ C.4 Detailed Illustration of Datasets
Appendix D Further Ablation Studies and Discussions
@@
Figure 16: TQA (LLM) vs. VTQA on spatial reasoning tasks. Each vertex on the spider plot represents the Avg Acc of a (LLM, VLM) pair with the same language model backbone. For LLM, we use the Text-only input (i.e., TQA); for VLM, we use Vision-text input (i.e., VTQA).
Impact of prompting techniques.
In line with our approach to sampling strategies, our primary goal in choosing prompting techniques is to report the best model performance given the same question. The prompt technique we use, which asks for step-by-step explanation is the most effective query among others in our initial studies. As a concrete example, we compare the original prompting strategy, "First, provide a concise answer in one sentence. Then, elaborate on the reasoning behind your answer in a detailed, step-by-step explanation" (step-by-step explanation), with a simpler prompt, Answer: (completion). Results in Table 4 show that the simpler completion prompt consistently underperforms compared to the step-by-step explanation prompt.
Input Modality Model Avg Acc (completion) Avg Acc (step-by-step explanation)
Text-only Mistral-7B 0.61 0.62
Vicuna-13B-1.5 0.25 0.46
Vicuna-7B-1.5 0.31 0.41
Vision-only LLaVA-1.6-Mistral-7B 0.47 0.47
LLaVA-1.6-Vicuna-13B 0.25 0.40
LLaVA-1.6-Vicuna-7B 0.26 0.31
Vision-text LLaVA-1.6-Mistral-7B 0.59 0.59
LLaVA-1.6-Vicuna-13B 0.26 0.48
LLaVA-1.6-Vicuna-7B 0.30 0.46
Table 4: Results for completion and step-by-step prompts.
Impact of temperature for decoding strategies.
We conducted an ablation study to examine how different temperatures affect the model performance. A higher temperature allows for more diversity in model responses. Most models consistently underperform when the temperature is set to 1.0 compared to 0.2 (our default) as shown in Table 5.
Input Modality Model Avg Acc (temperature=1) Avg Acc (temperature=0.2)
Text-only Mistral-7B 0.62 0.62
Vicuna-13B-1.5 0.37 0.46
Vicuna-7B-1.5 0.36 0.41
Vision-only LLaVA-1.6-Mistral-7B 0.37 0.47
LLaVA-1.6-Vicuna-13B 0.32 0.40
LLaVA-1.6-Vicuna-7B 0.26 0.31
Vision-text LLaVA-1.6-Mistral-7B 0.46 0.59
LLaVA-1.6-Vicuna-13B 0.37 0.48
LLaVA-1.6-Vicuna-7B 0.33 0.46
Table 5: Results by varying different temperatures for decoding strategies.
Appendix E Results on Spatial-Real task
Table 6 presents the performance on the Spatial-Real task. The same trends observed in synthetic tasks persist with real images (see VQA vs. VTQA, TQA (LLM) vs. VTQA, TQA (LLM) vs. VQA in Table 2). Notably, compared to synthetic tasks (Figure 5 and Figure 11), the overall accuracy on Spatial-Real is increased across all three input types (TQA, VQA, and VTQA). However, the modality gap (accuracy difference between VTQA and VQA) widens significantly, from 7.0% on synthetic tasks to 30.0% on Spatial-Real. This indicates that the performance disparity is more pronounced on natural images.
Input Format Model Average Accuracy
Text-only (TQA (LLM)) Vicuna-13B-1.5 0.845
Vicuna-7B-1.5 0.716
Mistral-7B 0.800
LLaMA-3-8B 0.884
Vision-only (VQA) CogVLM 0.419
Qwen-VL 0.329
LLaVA-1.6-Mistral-7B 0.368
CogAgent 0.400
LLaVA-1.6-Vicuna-7B 0.432
LLaVA-1.6-Vicuna-13B 0.445
Vision-text (VTQA) CogVLM 0.594
Qwen-VL 0.729
LLaVA-1.6-Mistral-7B 0.761
CogAgent 0.471
LLaVA-1.6-Vicuna-13B 0.832
LLaVA-1.6-Vicuna-7B 0.806
Table 6: Performance on the Spatial-Real task. The same trends still hold on real images: VQA vs. VTQA, TQA (LLM) vs. VTQA, and TQA (LLM) vs. VQA.
Appendix F Detailed Experimental Results
Results for proprietary models are summarized in Table 7. In Section 3 and Section 4, to clearly illustrate the impact of modality, we present the results averaged over all questions in Figure 6 for VQA vs. TQA and Figure 11 for VQA vs. VTQA. This section provides a comprehensive breakdown of results for individual questions. Detailed comparative results for Spatial-Map, Maze-Nav, and Spatial-Grid are shown in Table 8, Table 9, and Table 10, respectively. We compare LLM and VLM with the same language model backbone. For the VLM assessments, we consider inputs in both vision-only (VQA) and vision-text (VTQA) formats.
@@ Appendix F Detailed Experimental Results
Table 7: Detailed results for proprietary models.
@@
Table 8: Detailed results for the Spatial-Map task. We compare LLM and VLM with the same language model backbone. For VLMs, we consider both vision-only and vision-text input format. The averaged results are summarized in Figure 6 and Figure 11.
@@
Table 9: Detailed results for the Maze-Nav task. We compare LLM and VLM with the same language model backbone. For VLMs, we consider both vision-only and vision-text input format. The averaged results are summarized in Figure 6 and Figure 11.
@@
Table 10: Detailed results for the Spatial-Grid task. We compare LLM and VLM with the same language model backbone. The averaged results are summarized in Figure 6 and Figure 11.